y dx-4(x+y6)dy=0


1
Expert's answer
2021-07-16T02:20:32-0400

Solution

Writing DE in the form P(x,y)dx + Q(x,y)dy = 0 we’ll get P(x,y) = y, Q(x,y) = -4(x+y6)

∂P/∂y=1 ≠ ∂Q/∂x=-4

Let’s find function  M(x,y), called an integrating factor, for which equality ∂(M(x,y)P(x,y))/∂y=∂(M(x,y)Q(x,y))/∂x is satisfied.

-[∂P/∂y - ∂Q/∂x]/P=-5/y

So M(x,y) may be found from equation

dM/M = -[∂P/∂y - ∂Q/∂x]dy/P=-5dy/y

ln|M| = -5ln|y|+lnC

M = C/y5  

Multiplying DE by M(x,y) with arbitrary constant C=1

dx/y4-4(x/y5+y)dy = 0

Let’s find function u(x,y) for which du = uxdx+uydy = 0, where ux = 1/y4 and uy = -4(x/y5+y)

u(x,y) = ∫dx/y4 + f(y)  = x/y4 + f(y)  , where f(y) – arbitrary function.

uy = -4 x/y5 +f’(y) = -4(x/y5+y) => f’(y) = -4y => f(y) = -2y2 - C

Substituting this into expression for u(x,y) we’ll get

u(x,y) =  x/y4 - 2y2 - C

Therefore general solution of DE is x/y4 - 2y2 = C